# Low dose gamma irradiation study of ATLAS ITk MD8 diodes

### v1

@@ 1 Introduction
Previous studies of $\gamma$-irradiated $n^{+}$-in-$p$ diodes indicate that the bulk current increases linearly with total ionizing dose (TID), while the surface current saturates [3, 4]. Unfortunately, the minimum TID in these studies was 66 Mrad, making it impossible to determine at what TID the surface current saturates. Studies performed on $\gamma$-irradiated ATLAS ITk strip main sensors of the ATLAS18 type [5] and ITk strip modules using the equivalent sensors [6] showed a steep increase of the total current already at low TIDs of 11 krad. However, the ATLAS18 main sensors do not have a contactable guard ring, thus the surface and bulk current can not be measured separately.
@@ 2 Samples and irradiation
The study was carried out on two types of $n^{+}$-in-$p$ standard float zone diodes from the ATLAS18 ITk strip production wafers [7] manufactured by Hamamatsu Photonics K. K. [9] - regular MD8 diodes (0.8 cm $\times$ 0.8 cm) without a $p$-stop implant (MD8) and diodes with $p$-stop implant between the bias ring (BR) and the guard ring (GR), referred to as MD8p. Both diode types have an active area of 0.545 cm2 and a thickness of 320 µm. In the study, the MD8 diodes were part of a structure called “Mini&MD8”, which integrates a miniature strip sensor and an MD8 diode on a single silicon piece [8], see Figure 1. The MD8p diodes were diced individually.
@@
Figure 1: (a) The test structure “Mini&MD8” includes a miniature strip sensor, an MD8 diode (0.8 cm $\times$ 0.8 cm), and several other smaller diodes. (b) The diced MD8p diode with p-stop implant between BR and GR (0.8 cm $\times$ 0.8 cm).
@@ 3 Experimental methods and results
The tests were carried out using a probe station with controlled temperature (20$\pm$0.1)C and relative humidity (<1%) for both unirradiated and irradiated diodes. The irradiated diodes were tested before and after annealing, which was performed for 80 minutes at 60C.
@@ 3.1 Bulk, surface, and total currents measured before and after irradiation
Figure 2(a) displays total, bulk, and surface currents as a function of bias voltage measured both on MD8 and MD8p unirradiated diodes. Before irradiation, the bulk and surface currents are of a comparable magnitude and contribute approximately equally to the total current. The bulk and surface currents are slightly larger in MD8p diodes.
@@ 3.2 Annealing studies of bulk, surface, and total currents
The bulk current increases slightly with increasing time of annealing at 60C, see Figure 3(a). The surface current also increases with time; however, no annealing is observed at a bias voltage of 480 V (Figure 3(b)). It seems that annealing corrects the soft breakdown of the diode observed before annealing. The evolution of the total, bulk, and surface currents with annealing time measured at full depletion (at a bias voltage of 300 V) is displayed in Figure 4(a). The surface and bulk currents seem to saturate by 640 minutes of annealing at 60C.
The MD8p diode irradiated to TID of 15 krad initially underwent annealing for 80 minutes at 60C, followed by isochronal annealing with 20-minute steps and temperatures ranging from 80C to 300C. The surface I–V characteristics measured after each annealing step are shown in Figure 5(a). The annealing behavior of the total, bulk, and surface currents measured at a bias voltage of 300 V is displayed in Figure 5(b). As can be seen, the currents exhibit an increase up to 100C. However, at higher annealing temperatures the currents decrease significantly, returning to levels comparable to those before irradiation. This is demonstrated by Figure 4(b), which compares the I–Vs of the total, bulk, and surface currents before irradiation to those measured after irradiation and annealing at 300C. These results show that the radiation-induced defects caused by $\gamma$-irradiation are completely annealed at high temperatures.
@@ 3.3 Temperature dependence of total, bulk, and surface currents
$I(T)=AT^{2}\exp\left(-\frac{E}{2kT}\right),$ (3.1)
where $k$ is the Boltzmann constant ($k$ = 8.617 $\times$ 10-5 eV.K-1). Parameters $A$ and activation energy $E$ were treated as free parameters for total, bulk, and surface currents.
@@
Figure 6: (a) Surface currents of the MD8p diode irradiated to 100 krad measured at testing temperatures ranging from -50C to +20C, and plotted as a function of bias voltage. (b) Temperature dependence of total, bulk, and surface currents of the MD8p diode irradiated to 100 krad and biased to 150 V. Measured data is fitted by the formula 3.1..
It was assumed that $E$ might vary among $I_{\mathrm{BULK}}$, $I_{\mathrm{SURF}}$, and $I_{\mathrm{TOT}}$ due to a presence of the $\mathrm{SiO_{2}}$ layer on the surface of tested samples. Additionally, $E$ could vary with dopant concentration; however, no significant changes in radiation-induced defects in the silicon bulk were expected at the low $\gamma$-doses used in this study. The results show no differences in the activation energies obtained by fitting the temperature dependence of total, bulk, and surface currents by the formula 3.1. Additionally, the values of activation energies are consistent across all three samples irradiated by different TIDs. The fitted values of the activation energy, averaged over all three samples, are $E_{\mathrm{TOT}}=(1.198\pm 0.009)$ eV, $E_{\mathrm{BULK}}=(1.191\pm 0.012)$ eV, and $E_{\mathrm{SURF}}=(1.199\pm 0.009)$ eV, which are consistent with the Chiligarov’s result of (1.209±0.007)1.209$\pm$ 0.007)1.209 ± 0.007 ) eV [12].
@@ 4 Conclusions
This study investigates the effects of 60Co $\gamma$-rays on the total, bulk, and surface currents in $n^{+}$-in-$p$ MD8 diodes from ATLAS18 ITk production wafers, exposed to the TIDs ranging from 0.5 to 100 krad. Pre-irradiation, the magnitudes of the bulk and surface currents of MD8 diodes are comparable; however, post-irradiation, the increase in total current is primarily driven by the surface current, with bulk current remaining stable up to 100 krad. The surface current saturation was not observed up to the maximum TID of 100 krad used in this study, whereas in our previous study of $\gamma$-irradiated MD8 diodes [3], saturation of radiation-induced surface charges occurred at the minimal tested TID of 66 Mrad. The results thus clearly indicate that the saturation range lies between 100 krad and 66 Mrad. Additionally, the bulk and surface currents are slightly higher in MD8p diodes with $p$-stop implant between BR and GR than in the regular MD8 diodes without this $p$-stop implant.
Annealing experiments at 60C demonstrated that total, bulk, and surface currents exhibit slight increases with annealing time, especially at bias voltages close to full depletion. This current increase is also observed for temperatures up to 100C. However, for annealing temperatures above 100C, all three currents decrease significantly, returning to levels comparable to pre-irradiation values. Radiation-induced defects in both the bulk and surface, caused by $\gamma$-irradiation, appear to be completely annealed at high temperatures. For comparison, in reference [13], where $n^{+}$-in-$p$ diodes were studied after higher TIDs of $\gamma$-irradiation up to 100–200 Mrad, the current remains stable up to 200C, then steeply increases till the temperature of 260C, and subsequently decreases sharply to the maximum studied temperature of 300C. No differences were observed between the activation energies derived from temperature dependence measured for total, surface, and bulk currents.
@@ Acknowledgments
Copyright 2025 CERN for the benefit of the ATLAS Collaboration. CC-BY-4.0 license

### v2

@@ 1 Introduction
Silicon tracking detectors [1, 2] at the High-Luminosity LHC (HL-LHC) will be subjected to intense radiation, which induces electrically active defects in the silicon. Radiation damage arises primarily through two mechanisms: Non-Ionizing Energy Loss (NIEL) and Total Ionizing Dose (TID). Non-ionizing effects are referred to as “bulk damage” and are caused by displacement of crystal atoms which therefore leads to the generation of silicon lattice defects, the formation of point and cluster defects, and consequently an increase in deep-level trap states near the middle of the band gap [3]. In our study, silicon samples were irradiated with 60Co gamma rays, where displacement damage is caused by Compton electrons ($\leq 1$~ MeV) producing only point defects - since the energy is insufficient to create defect clusters. On a macroscopic scale, NIEL damages in solid state detectors cause an increase of leakage current (increase in noise), changes in material resistivity, the worsening of the charge collection properties due to trapping mechanism and eventually, the decrease of the carrier’s mobility and their lifetime. On the other hand, ionizing effects affect the insulating layer SiO2 and at the Si-SiO2 interface region. These effects are usually referred to as “surface damage” and are responsible for the build-up of trapped charge in the oxide, an increase in bulk oxide traps and interface traps [4] . In contrast to the silicon bulk, the creation of electron-hole pairs is not fully reversible in insulating materials such as the SiO2 passivation layer, where some electron-hole pairs become separated under an electric field. Due to very different mobility of electrons and holes most of the free electrons leave the SiO2, whereas the holes can drift toward the Si-SiO2 interface and become trapped, leading to an accumulation of positive oxide charges and interface states. The accumulation of positive charge attracts the negative charge from the bulk and the conductive electron channels are created. These effects lead to increased surface leakage currents and deteriorated sensor isolation. In earlier measurements on $p$-in-$n$ types of sensors it was observed that the density of oxide charges shows saturation at doses of the order of 100 krad [5]. The saturation value is technology-dependent and is larger if the detector is biased during irradiation.
Our previous studies of $\gamma$-irradiated $n^{+}$-in-$p$ diodes indicate that the bulk current increases linearly with TID, while the surface current saturates [6, 7]. Unfortunately, the minimum TID in these studies was 66 Mrad, making it impossible to determine at what TID the surface current saturates. Studies performed on $\gamma$-irradiated ATLAS ITk strip main sensors of the ATLAS18 type [8] and ITk strip modules using the equivalent sensors [9] showed a steep increase of the total current already at low TIDs of 11 krad. However, the ATLAS18 main sensors do not have a contactable guard ring, thus the surface and bulk current can not be measured separately.
@@ 2 Samples and irradiation
The study was carried out on two types of $n^{+}$-in-$p$ standard float zone diodes from the ATLAS18 ITk strip production wafers [10] manufactured by Hamamatsu Photonics K. K. [11]: regular MD8 diodes (0.8 cm $\times$ 0.8 cm) without a $p$-stop implant (MD8), and MD8p diodes, which include a $p$-stop implant between the bias ring (BR) and the guard ring (GR). Both diode types have an active area of 0.545 cm2 and a thickness of 320 µm. The additional p-stop implant ring (shown in red colour in Figure 1(b) isolates the currents that flow in the diode area from the edge region, especially of surface one, thus providing a better-defined diode volume and reducing edge current interference. The MD8p design allows for more precise studies of radiation-induced damage. In the study, the MD8 diodes were part of a structure called “Mini&MD8”, which integrates a miniature strip sensor and an MD8 diode on a single silicon piece [12], see Figure 1(a). The MD8p diodes were diced individually.
@@
Refer to caption   Refer to caption
Figure 1: (a) The test structure “Mini&MD8” includes a miniature strip sensor, an MD8 diode (0.8 cm $\times$ 0.8 cm), and several other smaller diodes. (b) The layout of MD8 and MD8p diodes, which is the same except for the extra p-stop implant between BR and GR ring in MD8p. [10]
@@ 3 Experimental methods and results
[…]
@@ 3.1 Bulk, surface, and total currents measured before and after irradiation
Figure 2(a) displays total, bulk, and surface currents as a function of bias voltage measured both on MD8 and MD8p unirradiated diodes. Before irradiation, the bulk and surface currents are of a comparable magnitude and contribute approximately equally to the total current. The bulk and surface currents are slightly larger in MD8p diodes. The cause of the difference is under investigation.
@@ 3.2 Annealing studies of bulk, surface, and total currents
The bulk current increases slightly with increasing annealing time at 60 C, as shown in Figure 3(a). The surface current also increases with time; however, at higher bias voltages (above $\sim 450\,\mathrm{V})$, the current increase becomes negligible, see Figure 3(b). These results suggest that annealing mitigates the soft breakdown of the diode observed before annealing at these bias voltages. The evolution of the total, bulk, and surface currents with annealing time measured at full depletion (at a bias voltage of 300 V) is displayed in Figure 4(a). The surface and bulk currents seem to saturate at about 640 minutes of annealing at 60 C.
The MD8p diode irradiated to TID of 15 krad initially underwent annealing for 80 minutes at 60 C, followed by isochronal annealing with 20-minute steps and temperatures ranging from 80 C to 300 C. The surface I–V characteristics measured after each annealing step are shown in Figure 5(a). The annealing behavior of the total, bulk, and surface currents measured at a bias voltage of 300 V is displayed in Figure 5(b). As can be seen, the currents exhibit an increase up to 100 C. However, at higher annealing temperatures the currents decrease significantly, returning to levels comparable to those before irradiation. This is demonstrated by Figure 4(b), which compares the I–Vs of the total, bulk, and surface currents before irradiation to those measured after irradiation and annealing at 300 C. These results show that the radiation-induced defects caused by $\gamma$-irradiation are completely annealed at high temperatures. For comparison, in reference [15], where $n^{+}$-in-$p$ diodes irradiated by $\gamma$-rays to TIDs in range 100 - 200 Mrad were studied, the leakage current remains stable up to 200C, then steeply increases till the temperature of 260C, and subsequently decreases sharply to the maximum studied temperature of 300C.
@@ 3.3 Temperature dependence of total, bulk, and surface currents
$I(T)=AT^{2}\exp\left(-\frac{E_{A}}{2kT}\right),$ (3.1)
where $k$ is the Boltzmann constant ($k$ = 8.617 $\times$ 10-5 eV.K-1). Parameters $A$ and activation energy $E_{A}$ were treated as free parameters for total, bulk, and surface currents.
@@
Figure 6: (a) Surface currents of the MD8p diode irradiated to 100 krad measured at testing temperatures ranging from -50 C to +20 C, and plotted as a function of bias voltage. (b) Temperature dependence of total, bulk, and surface currents of the MD8p diode irradiated to 100 krad and biased to 150 V. Measured data is fitted by the formula 3.1.
It was assumed that $E_{A}$ might vary among $I_{\mathrm{BULK}}$, $I_{\mathrm{SURF}}$, and $I_{\mathrm{TOT}}$ due to a presence of the $\mathrm{SiO_{2}}$ layer on the surface of tested samples. Additionally, $E_{A}$ could vary with dopant concentration; however, no significant changes in radiation-induced defects in the silicon bulk were expected at the low $\gamma$-doses used in this study. The results show no differences in the activation energies obtained by fitting the temperature dependence of total, bulk, and surface currents by the formula 3.1. Additionally, the activation energy values are consistent across all three samples irradiated with different TIDs, as shown in Table 1, where the fitted values of $E_{A}$ and $A$ for all three types of currents - total, bulk, and surface - are summarized for each tested diode. The fitted values of the activation energy, averaged over all three samples, are $E_{\mathrm{A(TOT)}}=(1.198\pm 0.009)$ eV, $E_{\mathrm{A(BULK)}}=(1.191\pm 0.012)$ eV, and $E_{\mathrm{A(SURF)}}=(1.199\pm 0.009)$ eV, which are consistent with the Chiligarov’s result of ($1.209\pm 0.007)$ eV [16].
TID $E_{\rm A(TOT)}$ $E_{\rm A(BULK)}$ $E_{\rm A(SURF)}$ $A_{\rm TOT}$ $A_{\rm BULK}$ $A_{\rm SURF}$
[krad] [eV] [eV] [eV] [nA/K2] [nA/K2] [nA/K2]
10 1.209 1.192 1.211 2.101 0.166 1.951
50 1.184 1.172 1.185 2.281 0.137 2.148
100 1.201 1.208 1.200 2.285 0.360 3.930
Table 1: Results of the fit: Activation energy $E_{A}$ and parameter AA for for total, surface, and bulk currents.
@@ 4 Conclusions
This study investigates the effects of 60Co $\gamma$-rays on total, bulk, and surface currents in $n^{+}$-in-$p$ MD8 diodes from ATLAS18 ITk production wafers, exposed to the TIDs ranging from 0.5 to 100 krad, with the aim of understanding current behavior during the initial operation of the new ATLAS tracker.
Before irradiation, the bulk and surface currents in MD8 diodes are comparable; after $\gamma$-ray exposure, the total current increases mainly due to a rise in surface current, with the bulk current staying constant up to maximum TID of 100 krad. No saturation of radiation-induced surface charges was observed within the studied range of TIDs, suggesting that the saturation threshold lies between 100 krad and the 66 Mrad value previously reported in Ref. [6].
Annealing at 60–100C causes a slight increase in bulk and surface currents, especially at bias voltages close to full depletion. In contrast, annealing above 100C significantly reduces both surface and bulk components, restoring current levels to pre-irradiation values. This suggests that defects induced by $\gamma$-irradiation in both the bulk and surface are effectively annealed at high temperatures.
No differences were observed between the activation energies derived from temperature dependence measured for total, surface, and bulk currents.
@@ Acknowledgments
Copyright 2025 CERN for the benefit of the ATLAS Collaboration. CC-BY-4.0 license.